# Atmospheric Characterisation with the Twinkle Space Telescope Following Advances from JWST Observations–B Atmospheric Characterisation with the Twinkle Space Telescope Following Advances from JWST Observations

### v1

@@ Abstract
The Twinkle Space Telescope is a satellite designed for spectroscopic observations of a wide range of extrasolar and solar system objects. Equipped with a 0.45 m diameter telescope and a spectrometer covering from 0.5 to 4.5 m simultaneously, Twinkle will be launched in a sun-synchronous, low-Earth orbit, and it is expected to operate for seven years. Twinkle is developed, managed and operated by Blue Skies Space (BSSL), a space science data company whose vision is to accelerate and expand the availability of new, high-quality datasets to researchers worldwide, complementing the space-observatories delivered by government space agencies.
Over its life-time, Twinkle will conduct large-scale survey programs. The scientific objectives and observational strategy of these surveys are defined by researchers who join the Science Team.
Leveraging advances made possible by recent observations with the James Webb Space Telescope, we present here updated simulations evaluating Twinkle’s observational capabilities in the context of exoplanet atmospheres. Through retrieval analyses of HD 209458 b, WASP-107 b, GJ 3470 b, and 55 Cnc e, we demonstrate how increasing observational investment enhances the retrieval of atmospheric parameters and molecular abundances. Our sensitivity study highlights Twinkle’s capability to detect less abundant/detectable molecules depending on the observing strategies adopted. This work provides practical guidance for developing targeted observational strategies to maximise Twinkle’s scientific return.
@@
The field of exoplanetary science has experienced unprecedented growth, with the number of confirmed exoplanets now exceeding 5,800. A significant proportion of these detections – approximately 4,000 – has been made using transit photometry from space-based surveys such as Kepler, K2, and TESS (Borucki et al., 2010; Howell et al., 2014; Ricker et al., 2014). Since the retirement of the Kepler/K2 mission in 2018, the TESS mission, operational since 2018, has significantly advanced planet discovery efforts. As of 12 March 2025, TESS has identified 7,525 TESS Objects of Interest (TOIs). After excluding false positives, 5,253 candidates remain confirmed by the TESS Follow-up Observing Program Working Group (TFOPWG) (Akeson & Christiansen, 2019). This wealth of discoveries of transiting planets around bright stellar hosts has been a game changer for atmospheric characterisation studies using transit, eclipse and phase-curve spectroscopy.
@@
The Chinese Space Station Telescope (Xuntian), scheduled for late 2026, will offer high-resolution multiband imaging and slitless spectroscopy across 0.255–1 m (Zhan, 2021).
The Twinkle Space Telescope, managed and operated by Blue Skies Space Ltd. (BSSL), will be dedicated to provide spectroscopic data for various science themes developed by its members (Jason et al., 2016; Savini et al., 2018; Edwards et al., 2019a, b; Archer et al., 2020; Stotesbury et al., 2022, 2024). Equipped with a 0.45-metre telescope and a spectrometer covering simultaneously from 0.5 to 4.5 m, Twinkle is capable of providing spectra for a wide range of targets such as brown dwarfs, stars, protoplanetary disks, Solar System objects and exoplanets. Two large survey programmes focusing on extrasolar targets and Solar System objects are being planned during the first three years of Twinkle’s expected seven years operations (Stotesbury et al., 2022). While BSSL is responsible for the satellite, the data analysis tools and facilitating the survey planning process, researchers who joins the survey program (Science Team) will decide the scientific objectives and observational strategy. Twinkle Science Team members have been tasked with delivering a comprehensive and scientifically optimised observation plan and list of targets.
@@
Previous studies by Edwards et al. (2019c) evaluated Twinkle’s expected performance to study the atmospheric composition and structure of three iconic exoplanets, HD 209458 b, GJ 3470 b, and 55 Cnc e. However, due to lack of IR spectroscopic data pre-JWST, those analyses were based on theoretical models and simplified assumptions. Recent JWST observations have revealed complexities in atmospheric chemistry, thermal and cloud structures that necessitate an update to these models (e.g., Xue et al., 2024; Beatty et al., 2024; Hu et al., 2024). More specifically, recent JWST observations of HD 209458 b by Xue et al. (2024) have provided a high-quality transmission spectrum from 2.3 to 5.1 m, revealing pronounced absorption features from CO2 and H2O. Based on thermochemical equilibrium assumptions, these results suggest a supersolar metallicity and a low C/O ratio (0.11). GJ 3470 b has emerged as a key target of interest due to its relatively transparent atmosphere and the detection of SO2 in its atmosphere, reinforcing the evidence of disequilibrium chemistry in lower-mass exoplanet atmospheres (Beatty et al., 2024). JWST’s eclipse observations of the super-Earth 55 Cnc e by Hu et al. (2024) have confirmed the presence of a volatile-rich atmosphere, potentially maintained by a magma ocean. In addition to these three exoplanets, we have added here the study of WASP-107 b: Welbanks et al. (2024) have identified the presence of several molecular species in the atmosphere of this planet, including H2O, CH4, CO, CO2, SO2, and NH3 when combining HST and JWST data.
| HD 209458 b | WASP-107 b | GJ 3470 b | 55 Cancri e | |||||||||
| Planet Type | Hot Jupiter | Warm Neptune | Warm Neptune | Super Earth | ||||||||
| Planet Mass () | ||||||||||||
| Planet Radius () | ||||||||||||
| Teq (K) | ||||||||||||
| Stellar Magnitude (V/K) |
| JWST Availability (m) | 2.3–4.5 | 2.4–4.5 | 2.45–4.97 | 3.94–12 | ||||||||
| HST Wavelength Range (m) | 0.5–0.9; 1.1–1.7 | 0.8–1.6 |
@@ 2.1 Twinkle’s Radiometric Model
Twinkle’s radiometric performance is assessed using the Twinkle Radiometric Tool (Stotesbury et al., 2022), adapted from the open-source radiometric software ExoRad (Mugnai et al., 2023). The radiometric model simulates observational and instrumental noise contributions, including photon noise, instrument emission, detector dark current, read noise, and zodiacal background. This model has been refined to incorporate recent advances in instrument design. It calculates realistic noise estimates, which in turn allows determination of the number of observations necessary to reach a specific SNR through observation stacking. The cumulative SNR achieved by combining multiple observations can be expressed as:
where is the number of observations, and is the signal-to-noise ratio (SNR) of a single observation. The signal is derived from the amplitude () of spectral features in transit spectra, approximated by:
@@
Previously, due to limited observational constraints, atmospheric thickness () was set as 1, 3, or 5 scale heights, with 5 scale heights being a commonly used assumption (e.g., Edwards et al., 2019c; Booth et al., 2024). However, with the aid of recent data release from JWST and its overlapping wavelength coverage with Twinkle, constraints on atmospheric composition have improved, enabling more precise estimates of atmospheric scale height and thickness above the optically thick continuum. Thus, we can derive a more realistic SNR and resolution needed for Twinkle to better retrieve the atmospheric chemical composition, and consequently, determine the number of observations required to achieve a given SNR. It is worth noting that the radiometric model employed in this study calculates Twinkle’s observational errors assuming complete transit/eclipse coverage. However, observational availability of targets are influenced by Earth obstruction due to Twinkle’s orbital configuration, causing Twinkle to sometimes capture partial transits or eclipse events, similar to the situation encountered by CHEOPS (Benz et al., 2021).
While for some exoplanets, the number of required observations is high, it is important to note that lowering spectral resolution can increase the SNR per wavelength bin and thus reduce the observational requirements. Although this method is widely used (e.g., Kreidberg et al., 2014; Stevenson et al., 2014), artificially reducing resolution in simulations introduces bias along the wavelength axis, particularly around spectral peaks. To prevent this bias, this study retrieves spectra at Twinkle’s full resolution. After the launch of Twinkle, the actual resolution adjustment will effectively improve this bias.
@@ 2.2 Twinkle’s Target Candidates
Twinkle will be placed in a 1200 km, sun-synchronous low-Earth orbit, with a 6am Local Time of Ascending Node (LTAN). The spacecraft’s observational capabilities are restricted by the Field of Regard (FoR) (a 40 degree cone around the anti-Sun vector), solar & lunar exclusion constraints and Earth obstruction across different seasons.(Stotesbury et al., 2022).
To systematically assess target observability, we employed the BSSL Twinkle Radiometric Tool to simulate observational noise for all confirmed exoplanets within Twinkle’s FOR possessing well-defined stellar and planetary parameters. Planetary and stellar data from the NASA Exoplanet Archive are taken as input into the radiometric tool to simulate noise levels for single transits or eclipses at Twinkle’s maximum resolving power. For TESS Objects of Interest (TOIs) lacking direct planetary mass measurements and thus not able to be processed with Radiometric Tool, masses were estimated using the methodology described by Chen & Kipping (2017). Approximately 200 TOIs within Twinkle’s FOR were subsequently excluded due to unavailable stellar mass data. Additionally, initial mass-radius relationships for gas giants between 8 and 22 Earth radii proved inconsistent with current empirical data. Hence, planetary masses for these TOIs were recalculated based on a log-normal distribution fitted to known transiting exoplanet populations, as detailed in Figure 15 and further described in Appendix B.
Our primary focus was on exoplanets with H2/He dominated atmospheres; thus, we assumed a mean molecular weight of 2.3 for initial assessments. Atmospheric thicknesses equivalent to 3 atmospheric scale heights were adopted uniformly for all candidate targets. Exoplanets predicted to achieve median SNR levels of 3, 5, or 7 within 30 transits or eclipses at Twinkle’s native spectral resolution were considered suitable candidates for detailed atmospheric characterisation.
@@ 2.3 Simulated Exoplanets and Spectral Retrievals
We employed the open-source retrieval code TauREx 3.2.2 (Al-Refaie et al., 2021) to simulate forward models for transmission and emission spectra and to perform spectral retrieval analyses. Stellar spectra for each target star were simulated using the PHOENIX stellar library (Husser et al., 2013). Chemical profiles were generated using the FastChem equilibrium chemistry plugin for equilibrium assumptions, supplemented by TauREx’s free chemistry feature for non-equilibrium chemistry scenarios. All atmospheric models assumed a plane-parallel geometry comprising 100 layers, including molecular absorption, Rayleigh scattering, grey clouds, and collision-induced absorption (CIA). For all exoplanets, CIA processes involving H2-H2 and H2-He were considered, with additional contributions from N2-N2 and O2-O2 specifically included for 55 Cnc e. Molecular opacity cross-sections were sourced from the ExoMol database (Tennyson et al., 2024), while CIA data were taken from the HITRAN database (Karman et al., 2019).
Retrieval analyses were performed using the Bayesian MultiNest algorithm (Feroz et al., 2009), employing 1000 live points and an evidence tolerance threshold of 0.5. Non-informative priors for molecular abundances were uniformly set on a log scale within the range of to , while other planetary parameters utilised priors constrained by previous literature of specific planets.
@@
For HD 209458 b, recent JWST observations by Xue et al. (2024) were combined with HST data from Saba et al. (2025). The simulated spectrum is fitted with observed data through minor adjustments. Similarly, to assess Twinkle’s capabilities for WASP-107 b, a simulated spectrum covering Twinkle’s wavelength coverage (0.5–4.5 m) was constructed using HST and JWST NIRSpec data as benchmarks. The atmospheric model included known molecular constituents identified by Sing et al. (2024), specifically H2O, CH4, NH3, CO, CO2, SO2, and H2S, alongside potential SiO detection proposed by Ma et al. (2025), and optical absorbers Na and K, with uniform abundances assumed.
@@
To evaluate the retrievability of minor atmospheric species whose spectral signatures are typically obscured by clouds or dominated by major molecules, we conducted a dedicated sensitivity study. Using WASP-107 b as the test case, we systematically amplified the abundances of minor species (CH4, SO2, NH3, and H2S) to identify the enhancement factor necessary for their spectral features to become clearly observable above simulated observational uncertainties. The baseline scenario employed the previously established retrieval setup for WASP-107 b, stacking six transits to achieve a median SNR of 10. Individual retrieval analyses were then performed separately for each molecule, progressively increasing their abundances, to quantify their retrieval thresholds under varying observational conditions.

### v2

# Atmospheric Characterisation with the Twinkle Space Telescope Following Advances from JWST Observations–D Atmospheric Characterisation with the Twinkle Space Telescope Following Advances from JWST Observations

@@ Abstract
The Twinkle Space Telescope is a satellite designed for spectroscopic observations of a wide range of extrasolar and solar system objects. Equipped with a 0.45 m diameter telescope and a spectrometer covering from 0.5 to 4.5 m simultaneously, Twinkle will be launched in a sun-synchronous, low-Earth orbit, and it is expected to operate for seven years. Twinkle is developed, managed and operated by Blue Skies Space (BSSL), a space science data company whose vision is to accelerate and expand the availability of new, high-quality datasets to researchers worldwide, complementing the space-observatories delivered by government space agencies. Over its life-time, Twinkle will conduct large-scale survey programs. The scientific objectives and observational strategy of these surveys are defined by researchers who join the Science Team. Leveraging advances made possible by recent observations with the James Webb Space Telescope, we present here updated simulations evaluating Twinkle’s observational capabilities in the context of exoplanet atmospheres. Through retrieval analyses of HD 209458 b, WASP-107 b, GJ 3470 b, and 55 Cnc e, we demonstrate how increasing observational investment enhances the retrieval of atmospheric parameters and molecular abundances. Our sensitivity study highlights Twinkle’s capability to detect less abundant/detectable molecules depending on the observing strategies adopted. This work provides practical guidance for developing targeted observational strategies to maximise Twinkle’s scientific return.
@@
The field of exoplanetary science has experienced unprecedented growth, with the number of confirmed exoplanets now exceeding 5,800. A significant proportion of these detections – approximately 4,000 – has been made using transit photometry from space-based surveys such as Kepler, K2, and TESS (Borucki et al., 2010; Howell et al., 2014; Ricker et al., 2014). Since the retirement of the Kepler/K2 mission in 2018, the TESS mission, operational since 2018, has significantly advanced planet discovery efforts. As of 12 March 2025, TESS has identified 7,525 TESS Objects of Interest (TOIs). After excluding false positives, 5,253 candidates remain confirmed by the TESS Follow-up Observing Program Working Group (TFOPWG) (Akeson and Christiansen, 2019). This wealth of discoveries of transiting planets around bright stellar hosts has been a game changer for atmospheric characterisation studies using transit, eclipse and phase-curve spectroscopy.
@@
The Twinkle Space Telescope, managed and operated by Blue Skies Space Ltd. (BSSL), will be dedicated to provide spectroscopic data for various science themes developed by its members (Jason et al., 2016; Savini et al., 2018; Edwards et al., 2019c, a; Archer et al., 2020; Stotesbury et al., 2022, 2024). Equipped with a 0.45-metre telescope and a spectrometer covering simultaneously from 0.5 to 4.5 m, Twinkle is capable of providing spectra for a wide range of targets such as brown dwarfs, stars, protoplanetary disks, Solar System objects and exoplanets. Two large survey programmes focusing on extrasolar targets and Solar System objects are being planned during the first three years of Twinkle’s expected seven years operations (Stotesbury et al., 2022). While BSSL is responsible for the satellite, the data analysis tools and facilitating the survey planning process, researchers who joins the survey program (Science Team) will decide the scientific objectives and observational strategy. Twinkle Science Team members have been tasked with delivering a comprehensive and scientifically optimised observation plan and list of targets.
@@
Previous studies by Edwards et al. (2019b) evaluated Twinkle’s expected performance to study the atmospheric composition and structure of three iconic exoplanets, HD 209458 b, GJ 3470 b, and 55 Cnc e. However, due to lack of IR spectroscopic data pre-JWST, those analyses were based on theoretical models and simplified assumptions. Recent JWST observations have revealed complexities in atmospheric chemistry, thermal and cloud structures that necessitate an update to these models (e.g., Xue et al., 2024; Beatty et al., 2024; Hu et al., 2024). More specifically, recent JWST observations of HD 209458 b by Xue et al. (2024) have provided a high-quality transmission spectrum from 2.3 to 5.1 m, revealing pronounced absorption features from CO2 and H2O. Based on thermochemical equilibrium assumptions, these results suggest a supersolar metallicity and a low C/O ratio (0.11). GJ 3470 b has emerged as a key target of interest due to its relatively transparent atmosphere and the detection of SO2 in its atmosphere, reinforcing the evidence of disequilibrium chemistry in lower-mass exoplanet atmospheres (Beatty et al., 2024). JWST’s eclipse observations of the super-Earth 55 Cnc e by Hu et al. (2024) have confirmed the presence of a volatile-rich atmosphere, potentially maintained by a magma ocean. In addition to these three exoplanets, we have added here the study of WASP-107 b: Welbanks et al. (2024) have identified the presence of several molecular species in the atmosphere of this planet, including H2O, CH4, CO, CO2, SO2, and NH3 when combining HST and JWST data.
| HD 209458 b | WASP-107 b | GJ 3470 b | 55 Cancri e | |
| Planet Type | Hot Jupiter | Warm Neptune | Warm Neptune | Super Earth |
| Planet Mass () | ||||
| Planet Radius () | ||||
| Teq (K) | ||||
| Stellar Magnitude (V/K) | 7.65 0.03 (V) 6.308 0.026 (K) | 11.592 0.208 (V) 8.637 0.023 (K) | 12.332 0.016 (V) 7.989 0.023 (K) | 5.95 0.023 (V) 4.015 0.036 (K) |
| JWST Availability (m) | 2.3–4.5 | 2.4–4.5 | 2.45–4.97 | 3.94–12 |
| HST Wavelength Range (m) | 0.5–0.9; 1.1–1.7 | 0.8–1.6 | 0.524–1.027; 1.1–1.7 | – |
@@ 2.1 Twinkle’s Radiometric Model
Twinkle’s radiometric performance is assessed using the Twinkle Radiometric Tool (Stotesbury et al., 2022), adapted from the open-source radiometric software ExoRad2 (Mugnai et al., 2023). The radiometric model simulates observational and instrumental noise contributions, including photon noise, instrument emission, detector dark current, read noise, and zodiacal background. This model has been refined to incorporate recent advances in instrument design (as elaborated in Stotesbury et al. (2024)). Detailed noise contributions for each target simulated in this study are presented in Appendix A. It is worth noting that the scope of this study covers only radiometric simulation, without detailed time-domain simulation, which will be addressed in a forthcoming paper. However, an average jitter value derived from ground tests is applied. The assumed instrument parameters are summarised in Table 9 of Appendix A.
The radiometric tool calculates realistic noise estimates, which in turn allows determination of the number of observations necessary to reach a specific SNR through observation stacking.
The cumulative SNR achieved by combining multiple observations can be expressed as:
where is the number of observations (each corresponding to a single transit or eclipse event), and is the signal-to-noise ratio (SNR) of a single observation. The signal is derived from the amplitude () of spectral features in transit spectra, approximated by:
@@
Previously, due to limited observational constraints, atmospheric thickness () was set as 1, 3, or 5 scale heights, with 5 scale heights being a commonly used assumption (e.g., Edwards et al., 2019b; Booth et al., 2024). However, with the aid of recent data release from JWST and its overlapping wavelength coverage with Twinkle, constraints on atmospheric composition have improved, enabling more precise estimates of atmospheric scale height and thickness above the optically thick continuum. Thus, we can derive a more realistic SNR and resolution needed for Twinkle to better retrieve the atmospheric chemical composition, and consequently, determine the number of observations required to achieve a given SNR. It is worth noting that the radiometric model employed in this study calculates Twinkle’s observational errors assuming complete transit/eclipse coverage. However, observational availability of targets are influenced by Earth obstruction due to Twinkle’s orbital configuration, causing Twinkle to sometimes capture partial transits or eclipse events, similar to the situation encountered by CHEOPS (Benz et al., 2021). The extent of transit or eclipse coverage depends on the target’s declination, the satellite’s orbital phase, and the ratio of the planet’s orbital period to Twinkle’s orbital period. Partial transits can still yield scientifically useful data but at reduced observing efficiency compared with full-coverage events. A detailed quantification of this efficiency factor, incorporating the effects of Earth obstruction and target-specific orbital geometry, will be addressed in a forthcoming time-domain simulation study.
While for some exoplanets, the number of required observations is high, it is important to note that lowering spectral resolution can increase the SNR per wavelength bin and thus achieve the desired SNR with fewer stacked observations. Although post-observation rebinning is widely used (e.g., Kreidberg et al., 2014; Stevenson et al., 2014), artificially reducing resolution in simulations introduces a loss of spectral information along the wavelength axis, particularly around spectral peaks. To avoid this, this study retrieves spectra at Twinkle’s full native resolution. After launch, rebinning applied to real observational data—rather than to simulated spectra—will mitigate this limitation, as it preserves the full information content of the original observations.
@@ 2.2 Twinkle’s Target Candidates
Twinkle will be placed in a 1200 km, sun-synchronous low-Earth orbit, with a 6am Local Time of Ascending Node (LTAN). The spacecraft’s observational capabilities are restricted by the Field of Regard (FoR) (a 40 degree cone around the anti-Sun vector), solar & lunar exclusion constraints and Earth obstruction across different seasons (Stotesbury et al., 2022).
To systematically assess target observability, we employed the BSSL Twinkle Radiometric Tool to simulate observational noise for all confirmed exoplanets within Twinkle’s FoR possessing well-defined stellar and planetary parameters. Planetary and stellar data from the NASA Exoplanet Archive are taken as input into the radiometric tool to simulate noise levels for single transits or eclipses at Twinkle’s maximum resolving power. For TESS Objects of Interest (TOIs) lacking direct planetary mass measurements and thus not able to be processed with Radiometric Tool, masses were estimated using the Forecaster tool (Chen and Kipping, 2017). Approximately 200 TOIs within Twinkle’s FoR were subsequently excluded due to unavailable stellar mass data. Additionally, planets with radii between 8 and 22 Earth radii fall outside the calibrated range of Forecaster, and the initial mass-radius relationships for gas giants proved inconsistent with current empirical data. Hence, planetary masses for these TOIs were recalculated based on a log-normal distribution fitted to known transiting exoplanet populations, as detailed in Figure 16 and further described in Appendix C. TOIs lacking sufficient stellar or planetary data for reliable mass estimation were excluded from the candidate list.
Our primary focus was on exoplanets with H2/He dominated atmospheres; thus, we assumed a mean molecular weight of 2.3 for initial assessments. Atmospheric thicknesses equivalent to 3 atmospheric scale heights were adopted uniformly for all candidate targets. Exoplanets predicted to achieve median SNR levels of 3, 5, or 7 within 30 transits or eclipses at Twinkle’s native spectral resolution were considered suitable candidates for detailed atmospheric characterisation. The threshold of 30 observations represents an assumed upper limit per target, determined by Twinkle’s total available observing time, the number of targets in the survey, and the accessibility constraints imposed by the satellite’s orbit over the mission lifetime. This choice of upper limit is an operational planning decision rather than a constraint imposed by Twinkle’s instrumental capabilities.
@@ 2.3 Simulated Exoplanets and Spectral Retrievals
We employed the open-source retrieval code TauREx 3.2.2 (Al-Refaie et al., 2021) to simulate forward models for transmission and emission spectra and to perform spectral retrieval analyses. Stellar spectra for each target star were simulated using the PHOENIX stellar library (Husser et al., 2013). Chemical profiles were generated using the FastChem equilibrium chemistry plugin (Stock et al., 2018) for equilibrium assumptions, supplemented by TauREx’s free chemistry retrieval mode. In this mode, the volume mixing ratio of each molecule is treated as a free parameter, constant with altitude, and retrieved independently without imposing chemical equilibrium constraints (Al-Refaie et al., 2021). While this approach does not explicitly model disequilibrium chemistry, it provides a model-agnostic framework that can capture departures from equilibrium, such as those arising from vertical quenching or photochemistry. All atmospheric models assumed a plane-parallel geometry comprising 100 layers, including molecular absorption, Rayleigh scattering, grey clouds, and collision-induced absorption (CIA). For all exoplanets, CIA processes involving H2-H2 and H2-He were considered, with additional contributions from N2-N2 and O2-O2 specifically included for 55 Cnc e. Molecular opacity cross-sections were sourced from the ExoMol database (Tennyson et al., 2024), while CIA data were taken from the HITRAN database (Karman et al., 2019).
Retrieval analyses were performed using the Bayesian MultiNest algorithm (Feroz et al., 2009), employing 1000 live points and an evidence tolerance threshold of 0.5. Table 2 summarises the prior distributions adopted for all retrieval parameters. Non-informative priors for molecular abundances were uniformly set on a log scale within the range of to , while other planetary parameters utilised priors constrained by previous literature of specific planets.
| Parameter | Prior type | Range |
| Common to all planets | ||
| All molecular log-VMRs | LogUniform | |
| HD 209458 b | ||
| (K) | Uniform | |
| factor | Uniform | |
| (Pa) | LogUniform | |
| WASP-107 b | ||
| (K) | Uniform | |
| factor | Uniform | |
| (Pa) | LogUniform | |
| GJ 3470 b | ||
| (K) | Uniform | |
| factor | Uniform | |
| (Pa) | LogUniform | |
| 55 Cnc e | ||
| (K) | Uniform | |
| (K) | Uniform | |
| factor | Uniform | |
| (Pa) | LogUniform | |
| (Pa) | LogUniform | |
To quantify the detectability of individual molecular species, we employ the Likelihood Ratio Test (LRT) following the methodology described by Tessenyi et al. (2013). To guard against unreliable MAP estimates for poorly constrained parameters, we perform the LRT using both the MAP and posterior median values; a molecule is classified as a “Detection” only if both tests independently exceed the threshold. The full LRT procedure is detailed in Appendix D. The more conservative (lower) detection significance is reported in the results tables; molecules below the threshold are classified as “Upper Limit”.
@@
For HD 209458 b, recent JWST observations by Xue et al. (2024) were combined with HST data from Saba et al. (2025). The simulated spectrum was aligned with the observed data by applying a vertical offset to reconcile differences in absolute transit depth calibration between the HST and JWST instruments at overlapping wavelengths. Molecular abundances and temperature profile parameters were then fine-tuned to optimise the fit to the combined HST+JWST dataset. Similarly, to assess Twinkle’s capabilities for WASP-107 b, a simulated spectrum covering Twinkle’s wavelength coverage (0.5–4.5 m) was constructed using HST and JWST NIRSpec data as benchmarks. The atmospheric model included known molecular constituents identified by Sing et al. (2024), specifically H2O, CH4, NH3, CO, CO2, SO2, and H2S, alongside potential SiO detection proposed by Ma et al. (2025), and optical absorbers Na and K, with uniform abundances assumed.
@@
To evaluate the retrievability of minor atmospheric species whose spectral signatures are typically obscured by clouds or dominated by major molecules, we conducted a dedicated sensitivity study. Using WASP-107 b as the test case, we systematically amplified the abundances of minor species (CH4, SO2, NH3, and H2S) to identify the enhancement factor necessary for their spectral features to become clearly observable above simulated observational uncertainties. The baseline scenario employed the previously established retrieval setup for WASP-107 b, stacking six transits to achieve a median SNR of 10. For each molecule, the abundance was scaled from its baseline value by amplification factors ranging from 5 to 80 (the specific factors tested for each species are listed in Table 8), while all other atmospheric parameters were held fixed at their baseline values. Independent retrieval analyses were then performed at both SNR = 5 and SNR = 10 to determine the minimum amplification factor required for a constrained retrieval of each species.